The area of a square is increasing at the rate of 28cm2 s-1. Find the increasing rate of the length of a side,x when the area of the square is 49cm2 .


1
Expert's answer
2022-02-08T08:16:35-0500

Solution;

If the side of a square x;

Area,"A=x^2"

"\frac{dA}{dt}=2x\frac{dx}{dt}"

Make "\frac{dx}{dt}" subject of the formula;

"\frac{dx}{dt}=\frac{1}{2x}\frac{dA}{dt}"

But;

"A=x^2=49" therefore; "x=7cm"

"\frac{dA}{dt}=28cm^2s^{-1}"

"\frac{dx}{dt}=\frac{1}{2×7cm}(28cm^2s^{-1})"

"\frac{dx}{dt}=2cms^{-1}"

The rate of increase of the length is 2cm/s.